the following costs were incurred by odd a mfg cost

DM =105200

DL = 221400

MOH = 98600

based on the above data

a.determine prime cost

b.conversion cost

c.mfg cost


1
Expert's answer
2021-08-23T10:42:31-0400

a.  prime cost=DM+DL=105 200+221 400=326 600

b. conversion cost=DL+MOH=221 400+98 600=320 000

c. mfg cost=DM+DL+MOH=105 200+221 400+98 600=425 200